OpenStudy (anonymous):

If y = -4.8 when x = -1.6, find x when y = -24. Determine the Solution.

OpenStudy (anonymous):

y=-4.8 y=3 x=-1.6 x=1/3y y=-24 x=1/3(-24) x=-8